OpenStudy (anonymous):

Is it true that 2/3 = 0.67 ? Explain your answer

OpenStudy (accessdenied):

2/3 is not quite 0.67. It is close, but they are not the same. If you wanted to see for yourself, you could multiply both sides by 3: 2/3 = 0.67 3*2/3 = 3*0.67 Multiplying and dividing by 3 cancels each other, and the right side we can multiply out. 2 = 2.01 Close! But not exact!

OpenStudy (mathmath333):

do it and see in calculator

Nnesha (nnesha):

i mena when we divide 2/3 we get 0.66666666666666666 so we can say that its .67 isn't that right ??

OpenStudy (accessdenied):

0.67 is just an approximation of 2/3, really. 2/3 as a decimal is an infinite line of 6: 0.6666...., but we can say 2/3 is approximately 0.67, or approximately 0.6667, or 0.666666667. Each one is just a little closer.

Nnesha (nnesha):

and when woe multiply by 3 so its 2 =2.01 when we do nearest whole number so its 2 ???right

Nnesha (nnesha):

ohhh ok sir i trust you @accessDenied @JerryCharles sorry

OpenStudy (accessdenied):

No worries! It is natural to want to shorten it by rounding since writing an infinite number of 6 is time consuming. Just be aware that any rounding will change the true value slightly as an expense to having a more compact number. The way I'd write it is with two "squiggly" lines: \( 2/3 \approx 0.67 \), meaning its "close to" 0.67.

Nnesha (nnesha):

yeah right :)

Nnesha (nnesha):

@AccessDenied explained already

OpenStudy (accessdenied):

If 2/3 = 0.67 is a confusing example, it is sort of like saying 0 = 0.01. They're close, but they are not equal. Can you see that a bit more clearly? Or, in terms of scaling up, you could technically round 56 to 100 (to the nearest hundreds place). 56 = 100 is pretty hard to say!

OpenStudy (accessdenied):

So 2/3 = 0.67 is false. They're close numbers, but they are not -exactly- equal like the equals sign implies.

OpenStudy (accessdenied):

That may work. The argument of multiplying both sides by 3 to get different numbers is very strong as well: 2/3 = 0.67 2/3 * 3 = 0.67 * 3 2 = 2.01 2 is not equal to 2.01, so 2/3 is not equal to 0.67 Whatever you think your teacher will consider a good enough explanation!

OpenStudy (accessdenied):

Glad to help! :)